Five times the sum of the digits of a two-digit number is 13 less than the original number. If you reverse the digits in the two-digit number, four times the sum of its two digits is 21 less than the reversed two-digit number.
The difference of the original two-digit number and the number with reversed digits is

Answer by amarjeeth123(569)

Let the tens digit be x and the units digit be y.
5(x+y)=10x+y-13
5x=4y+13.......equation 1
4(x+y)=10y+x-21
3x=6y-21.......equation 2
x=2y-7.........equation 2
Substituting 2 in 1 we get,
5(2y-7)=4y+13
6y=48
y=8
x=16-7=9
The difference of the original two-digit number and the number with reversed digits is 98-89 i.e.9